Find the number of grams of 1 equiy EDA you need in a mixture of 1mL of 5 HM RuMpIXDME molar mass of 693.81) solution in Benzene
Answer #1

1ml of 5\muM of RuMpIXDME contain = 5 ×10-6 ×1/1000 =5×10 -9 moles of RuMpIXDME

For equivalent, we have equimolar solution of EDA and RuMpIXDME

So moles of EDA = 5× 10-9 moles

You have provided molar mass or full name of EDA.

If you multiple molar mass of EDA with 5× 10-9, then you have number of grams of EDA.

Here you need molar mass of EDA not molar mass of RuMpIXDME.